OpenStudy (anonymous):

Stan's statement: Equation - 5+2(x+4)=10-5(x-3) Step 1 - 5+2x+8=10-5x+15 Step 2 - 13+2x=25-5x Step 3 - 13-3x=25 Step 4 - -3x=12 Step 5 - -3x/-3=12/-3 Step 6 - x=-4 Which statement about Stan's solution is true? A. Stan's solution is correct. B. Stan has a mistake in Step 1. C. Stan has a mistake in Step 2. D. Stan has a mistake in Step 3. My work: Equation - 5+2(x+4)=10-5(x-3) 5+2x+8=10+-5x+15 13+2x=25+-5x +5x= +5x 13+7x=25 -13 =-13 7x/7=12/7 x=1 7/10 B. Stan has a mistake in Step 1. Is this correct? Did I make a mistake in my work?

Directrix (directrix):

Are you using - to mean minus or just a dash? In Equation - 5+2(x+4)=10-5(x-3), is the equation supposed to be - 5+2(x+4)=10-5(x-3) OR 5+2(x+4)=10-5(x-3)? It is dangerous to use dashes in mathematics.

Directrix (directrix):

I thought that Stan's equation is this: - 5+2(x+4)=10-5(x-3). I got the same thing he did for his first step.

Directrix (directrix):

In your work, I think you lost the negative sign in front of the 5. My work: Equation - 5+2(x+4)=10-5(x-3) 5+2x+8=10+-5+15

OpenStudy (anonymous):

Well, I don't usually use dashes, it is just hard for me to type math problems on a computer. The equation is 5+2(x+4)=10-5(x-3).

Directrix (directrix):

I got this for the first step which is what Stan got: 5+2x+8=10-5x+15 5+2x+8=10+-5+15 yours You left out the x on +-5 + 15 but I think you got what Stan did.

Directrix (directrix):

Do you think that this step Stan did is correct: Step 3) 13-3x=25

Directrix (directrix):

Let's look at it: Step 2) 13 +2x = 25 - 5x +5x = + 5x ---------------------------- 13 + 7x = 25

Directrix (directrix):

7x/7=12/7 x=1 7/10 12/7 = 1 5/7

Directrix (directrix):

Step 3 is where I think Stan made an error. What do you think: Step 3: 13-3x=25

Directrix (directrix):

Because: Step 2) 13+2x=25-5x 13 + 5x + 2x = 25 Step 3) 13+ 7x = 25 That is not what Stan got for Step 3.

OpenStudy (anonymous):

No, Stan messed up Step 3. Step 1 and 2 in my work looked the same as Stan's. Stan subtracted 5 from 2x instead of adding 5 to 2x.

OpenStudy (anonymous):

Thank you.

Directrix (directrix):

You are welcome.